(a) Two microwave frequencies are authorized for use in microwave ovens: 900 and 2560 MHz. Calculate the wavelength of each.

(b) Which frequency would produce smaller hot spots in foods due to interference effects?


1
Expert's answer
2022-01-24T11:26:23-0500

Wavelength

"\lambda_1=\frac{c}{\nu_1}"

"\lambda_1=\frac{3\times10^8}{900\times10^6}=0.33m"

"\lambda_2=\frac{c}{\nu_2}"

"\lambda_2=\frac{3\times10^8}{2560\times10^6}"

"\lambda_2=0.1171m"

(B)

2560 MHz( has the smallest wavelength)

"\lambda=0.1171m"

This is because a small wavelength microwave gives a pattern with which is interference with waves and have a closer hotspot